OpenStudy (anonymous):

rulnick one quick question about 300(A+12) = 1000A 3A+36=10A

OpenStudy (anonymous):

how did you get the 3A?

OpenStudy (rulnick):

300(A+12) = 1000A distribute the 300 ... 300A + 3600 = 1000A divide both sides by 100 ... 3A+36=10A

OpenStudy (anonymous):

i see thanks